### v2

# Private Continual Counting of Unbounded Streams

Abstract
We study the problem of differentially private continual counting in the unbounded setting where the input size $n$ is not known in advance. Current state-of-the-art algorithms based on optimal instantiations of the matrix mechanism [li_matrix_nodate] cannot be directly applied here because their privacy guarantees only hold when key parameters are tuned to $n$. Using the common ‘doubling trick’ avoids knowledge of $n$ but leads to suboptimal and non-smooth error. We solve this problem by introducing novel matrix factorizations based on logarithmic perturbations of the function $\frac{1}{\sqrt{1-z}}$ studied in prior works, which may be of independent interest. The resulting algorithm has smooth error, and for any $\alpha>0$ and $t\leq n$ it is able to privately estimate the sum of the first $t$ data points with $O(\log^{2+2\alpha}(t))$ variance. It requires $O(t)$ space and amortized $O(\log t)$ time per round, compared to $O(\log(n)\log(t))$ variance, $O(n)$ space and $O(n\log n)$ pre-processing time for the nearly-optimal bounded-input algorithm of henzinger_almost_2023. Empirically, we find that our algorithm’s performance is also comparable to theirs in absolute terms: our variance is less than $1.5\times$ theirs for $t$ as large as $2^{24}$.
1 Introduction
Differentially private counting under continual observation [chan2011private, dwork_differential_nodate] refers to the problem of maintaining accurate running totals over streams of sensitive data. It has attracted a great deal of attention in recent years [andersson_smooth_nodate, andersson_streaming_2024, andersson_continual_nodate, bienstock_dmm_2024, denisov2022improved, dvijotham_efficient_2024, fichtenberger2023constant, henzinger_almost_2023, henzinger2024unifying, mcmahan_inversion_2025] for its wide-ranging applications in optimization [choquette2023amplified, choquette2022multi, denisov2022improved, dwork2014analyze, kairouz2021practical, mcmahan2024hassle], as well as online learning and the private analysis of streaming data more broadly [agarwal2017price, cardoso2022differentially, jain_differentially_nodate, smith_nearly_nodate, upadhyay2019sublinear]. It has also been used in many large-scale deployments of differential privacy, such as Google’s private next-word prediction [mcmahan2022federated] and Smart Select models [hartmann2023distributed]. To a large extent, this flurry of activity has been prompted by recent algorithmic improvements in matrix factorizations for streaming data [denisov2022improved, fichtenberger2023constant, henzinger_almost_2023, li_matrix_nodate], which have dramatically improved privacy/utility tradeoffs compared to classical approaches based on binary trees [chan2011private, dwork_differential_nodate, honaker2015efficient].
One challenge when applying matrix-based algorithms to streaming settings is that they largely assume access to the input size $n$ as a parameter. While this is a very natural assumption in contexts like private model training, researchers from the beginning have also motivated their work with appeals to applications like public health [dwork_differential_nodate] where it is much less clear how the assumption could hold. In fact, it is known that difficulties related to unbounded inputs have influenced practical deployments of DP [amin2024practical]: famously, after Apple announced that it planned to use differential privacy to collect user data, tang2017privacy discovered that their implementation achieved this by guaranteeing privacy only over a single day while allowing privacy loss per-user to grow without bound over time.
Simple solutions to this problem all carry fundamental tradeoffs. Lifting a bounded-input algorithm to the unbounded case with some form of doubling trick [chan2011private] asymptotically preserves performance, but it also leads to non-smooth growth in error over time, undermining one of the major selling points of matrix methods over binary trees [henzinger_almost_2023]. Using a bounded matrix mechanism alongside the classical query model [dwork2004privacy, dwork2006calibrating] and refusing to produce additional outputs once our privacy budget is exhausted preserves performance and smoothness, but is not truly unbounded and is unlikely to be satisfying in practice. Finally, simple algorithms like adding independent noise to each data point are trivially unbounded with smooth noise, but have extremely sub-optimal error. Our main contribution in this paper is a matrix-based algorithm that overcomes this trilemma: to the best of our knowledge, it is the first private counting algorithm that is simultaneously smooth and unbounded with almost-optimal asymptotic error. It is also practical to implement, with essentially the same complexity as the foundational algorithm of henzinger_almost_2023, and we show empirically that it enjoys small error for realistic input sizes and not just asymptotically.
2 Background and Related Work
2.1 Background
Continual (Binary) Counting.
In continual counting [dwork_differential_nodate, chan2011private], the input consists of a (potentially unbounded) stream of data $x_{1},x_{2},\ldots$. At each time step $t$, the algorithm receives $x_{t}$ and then produces an estimate of the partial sum $S_{t}=\sum_{i=1}^{t}x_{t}$. For the sake of simplicity, we will assume that $x_{t}$ is a single bit, but our results are easy to generalize to $x_{t}\in\mathbb{R}^{d}$.
Differential Privacy under Continual Observation.
Define two (potentially unbounded) streams $x_{1},x_{2},\ldots$ and $x^{\prime}_{1},x^{\prime}_{2},\ldots$ to be neighboring if there is at most one time step $t$ where $x_{t}\neq x^{\prime}_{t}$. Let $X_{n}$ (resp. $X^{\prime}_{n}$) denote the first $n$ data points in each stream. We say that a mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta_{priv})$-differential privacy (DP) [dwork2014algorithmic] in the continual observation model [dwork_differential_nodate] if for all neighboring streams $X,X^{\prime}$, all $n\in\mathbb{Z}^{+}$, and all measurable events $S$, we have $\mathbb{P}(\mathcal{M}(X_{n})\in S)\leq e^{\varepsilon}\mathbb{P}(\mathcal{M}(X^{\prime}_{n}))+\delta_{priv}$. Here, $\mathcal{M}(X)$ represents the full sequence of outputs produced by $\mathcal{M}$ on input $X$. This reduces to the standard definition when the streams are finite; the additional quantification over $n$ is necessary to make sense of the definition for unbounded streams. The mechanisms we study all use additive Gaussian noise, and so can readily be shown to satisfy other notions of indistinguishability as well, such as Rényi DP mironov2017renyi, zCDP bun2016concentrated, or Gaussian DP dong2022gaussian. By virtue of using Gaussian noise, they also automatically satisfy DP in the stronger adaptive setting where $x_{t}$ is permitted to depend on earlier outputs, which is important for learning applications  [denisov2022improved].
Two measures of utility are standard in the literature: the root mean squared additive error and the expected maximum additive error, defined respectively as the maximums over $X_{n}$ of:
$\operatorname{err}^{\ell_{2}}_{n}(\mathcal{M}):=\sqrt{\frac{1}{n}\mathbb{E}_{\mathcal{M}}\big(\sum_{t\leq n}(\lVert S_{t}-\mathcal{M}(X_{t})\rVert_{2}^{2}\big)}$ $\operatorname{err}^{\ell_{\infty}}_{n}(\mathcal{M}):=\mathbb{E}_{\mathcal{M}}\big(\max_{t\leq n}~\lVert S_{t}-\mathcal{M}(X_{t})\rVert_{\infty}\big)$
When $\mathcal{M}$ uses additive Gaussian noise, both expectations and matching high-probability bounds can be derived from $\mathbb{V}(\mathcal{M}(X_{t}))$, the (non-asymptotic) variance of the noise added as a function of time. This is the main metric we will report when comparing the performance of different algorithms. We additionally say that a mechanism has smooth error if $\mathbb{V}(\mathcal{M}(X_{t}))$ is a smooth function of $t$. This is a desirable property because $\operatorname{err}^{\ell_{\infty}}_{n}$ depends on the maximum variance across all time steps, which can be significantly higher than the average variance when error is non-smooth [andersson_smooth_nodate].
2.2 Related Works
Differential privacy under continual observation was introduced by dwork_differential_nodate and chan2011private, who independently proposed the binary mechanism. Conceptually, this algorithm computes partial sums with the help of a binary tree where leaf nodes contain private estimates of individual data points, and internal nodes contain independent estimates of the subsequences spanned by their children. The binary mechanism of dwork_differential_nodate used Laplacian noise to satisfy $(\varepsilon,0)$-DP with $\operatorname{err}_{n}^{\ell^{\infty}}=O(\log^{2.5}(n))$, which was improved to $O(\log^{2.5}(t))$ by chan2011private. This was later reduced to $O(\log(t)\sqrt{\log n})$ by jain_differentially_nodate by using Gaussian noise instead. A straightforward optimization of the algorithm follows from the observation that at a given time step, it is only necessary to store $\log_{2}(n)$ nodes of the tree in memory. Subsequent work by honaker2015efficient additionally showed that it is possible to reduce the error by roughly a factor of 2 by making efficient use of multiple independent estimates for the same partial sum.
Both the original binary mechanism and Honaker’s variant have very uneven noise scale across time steps. Recently, andersson_smooth_nodate proposed a smooth variant of the binary mechanism which equalizes the noise distribution across time steps, further improving $\operatorname{err}^{\ell_{\infty}}_{n}$ by a constant factor.
The binary mechanism can be seen as a special case of the general matrix method  [li_matrix_nodate]. Given a dataset $x$ and a matrix $A=LR$, the matrix method estimates $Ax$ by $L(Rx+z)=Ax+Lz$. This satisfies DP when $z$ is a Gaussian scaled to $\rVert R\rVert_{1\to 2}$, the maximum $\ell_{2}$ norm over the columns of $R$. The final error also scales with $\lVert L\rVert_{2\to\infty}$, the maximum $\ell_{2}$ norm over the rows of $L$, and so by choosing $L$ and $R$ carefully we can achieve error depending on the factorization norm $\gamma_{2}(A)=\min\{\lVert L\rVert_{2\to\infty}\lVert R\rVert_{1\to 2}\mid LR=A\}$. This is often much more accurate than calibrating $z$ to $A$ directly. In the case of continual counting, $A=M_{\textit{count}}$, the all-ones lower-triangular matrix.
This matrix-based approach has rapidly become the standard in the field. edmonds_power_2019 proved a lower bound of $\Omega(\log n)$ on the per-time-step $\ell_{\infty}$ error of continual counting by lower-bounding $\gamma_{2}(M_{\textit{count}})$. denisov2022improved approached the problem of finding optimal decompositions as a convex optimization problem, building on earlier work in the offline setting  [yuan2016optimal]. They proved that, for lower-triangular matrices like $M_{\textit{count}}$, it is sufficient to consider only decompositions where $L$ and $R$ are both lower-triangular. This is a particularly nice result because it implies there is no extra cost for using the matrix method in streaming settings. Subsequently, fichtenberger2023constant provided an explicit factorization $L=R=M_{\textit{count}}^{1/2}$ and studied its $\operatorname{err}^{\ell_{\infty}}_{n}$ error, and henzinger_almost_2023 proved that the same decomposition is simultaneously nearly optimal in terms of $\operatorname{err}^{\ell_{2}}_{n}$. More recent work has investigated various computational modifications of this factorization, trading a small amount of accuracy for efficiency by approximating the optimal decomposition in more compact ways, e.g. through finite recurrences [dvijotham_efficient_2024, andersson_space_2025] or by binning together similar values [andersson_streaming_2024].
A distinct line of work has investigated the unbounded case. In their seminal work, chan2011private proposed a ‘hybrid’ mechanism which uses a variant of the classic doubling trick to match the asymptotic $\operatorname{err}_{n}^{\ell_{\infty}}$ rate of the binary mechanism when $n$ is unknown. Their technique is generic and can be used to ‘lift’ any fixed-size algorithm into one that works on unbounded streams with the help of a (potentially inefficient) unbounded algorithm — in section 5 we investigate a version of this mechanism instantiated with Algorithm 1. An alternative strategy for handling unbounded inputs is to only permit queries based on sliding windows or decaying sums [bolot2013private, upadhyay2019sublinear]. The dual approach, proposed by bolot2013private and recently extended by andersson_continual_nodate, instead relaxes the classical privacy definition to allow DP guarantees for past data points to gradually expire over time.
Our work is most directly inspired by the rational approximation methods of dvijotham_efficient_2024, and in particular by their use of the equivalence between certain matrix factorizations and generating functions. While they search for simple approximations of functions to derive computationally efficient factorizations, we take the opposite approach by searching for more complicated functions to derive matrix factorizations that are usable in the unbounded setting. Because we are not directly concerned with reducing computational complexity below $O(n)$ space and $O(n\log n)$ time, we will generally take the $M_{\textit{count}}^{1/2}$ mechanism of henzinger_almost_2023 and fichtenberger2023constant (“Sqrt Matrix” in Table 1 and Figure 2) as the baseline against which our algorithm is compared.
Table 1: Asymptotic Comparison of Private Counting Algorithms
Mechanism $\mathbb{V}(\mathcal{M}(X_{t}))\cdot\frac{\varepsilon^{2}}{\log(1/\delta_{priv})}$ Time for $t$ outputs Space Smooth Unbounded
Binary  [chan2011private, jain_differentially_nodate]* $O(\log(t)\log(n))$ $O(t)$ $O(\log t)$ No No
Hybrid Binary  [chan2011private]* $O(\log^{2}t+\log t)$ $O(t)$ $O(\log t)$ No Yes
Smooth Binary [andersson_smooth_nodate] $O((\log n+\log\log n)^{2})$ $O(t)$ $O(\log t)$ Yes No
Sqrt Matrix  [henzinger_almost_2023, fichtenberger2023constant] $O(\log(t)\log(n))$ $O(n\log n)$ $O(n)$ Yes No
Algorithm 1 $O(\log^{2+2\alpha}t)$ $O(t\log t)$ $O(t)$ Yes Yes
Hybrid Matrix $O(\log^{2}t+\log^{2+2\alpha}\log t)$ $O(t\log t)$ $O(t)$ No Yes
(*) Originally satisfied $\varepsilon$-DP with Laplacian noise; numbers reported are for the Gaussian variant. A mechanism is smooth if its variance is a smooth function of $t$, and it is unbounded if it doesn’t require $n$ as input. Algorithms with bolded names are original to this work. Constant factors differ significantly between algorithms — we investigate this more closely in section 5 and Figure 2.
3 Overview of Results
3.1 Preliminaries and Notation
The lower-triangular Toeplitz matrix corresponding to a sequence $a_{0},a_{1},\ldots$ is defined as the square matrix whose $i,j$ entry is $a_{i-j}$ when $i\geq j$ and 0 otherwise. For a given sequence, we will denote the corresponding matrix as $\operatorname{LTToep}(a_{0},a_{1},\ldots)$. All matrices we consider will be infinite, but for practical computations we will work with $A]_{n}$, the leading submatrix formed by the first $n$ rows and columns of $A$. The special matrix $M_{\textit{count}}$ is defined as $\operatorname{LTToep}(1,1,1,\ldots)$.
$\operatorname{LTToep}$ matrices are closed under multiplication, and $\operatorname{LTToep}(S)\operatorname{LTToep}(S^{\prime})=\operatorname{LTToep}(S\ast S^{\prime})$, where $S\ast S^{\prime}$ denotes the convolution of the sequences $S$ and $S^{\prime}$. This means that multiplication of $\operatorname{LTToep}$ matrices is commutative and can be performed in $O(n\log n)$ time using the Fast Fourier Transform (FFT). Given an analytic function $f(z)$ with Taylor coefficients $a_{n}=[z^{n}]f(z)$, we will abuse notation and write $\operatorname{LTToep}(f)$ for $\operatorname{LTToep}(a_{0},a_{1},\ldots)$. With this notation, applying the Cauchy product to their Taylor series gives us that $\operatorname{LTToep}(f)\operatorname{LTToep}(g)=\operatorname{LTToep}(fg)$. We refer the curious reader to the full version of dvijotham_efficient_2024 for an excellent and detailed introduction to $\operatorname{LTToep}$ matrices in this context, including proofs of the properties just described.
Finally, we note that we reserve the symbol $\delta$ to refer to the exponent of the $\log\log$ term in Equation 1. When we need to refer to the privacy parameter of a DP algorithm, we will use $\delta_{priv}$.
3.2 Main Result
Our main result is the following theorem and its associated algorithm (Algorithm 1):
Theorem 1.
For all $\alpha>0$, there exists an infinite lower-triangular Toeplitz matrix factorization $L,R\in\mathbb{R}^{\infty\times\infty}$ with the following properties:
  • Joint Validity: For all $n\in\mathbb{Z}^{+}$, $L]_{n}[R]_{n}=[M_{\textit{count}}]_{n}$
  • Bounded Sensitivity: $\lim_{n\to\infty}\lVert[R]_{n}\rVert_{1\to 2}=C<\infty$ for some computable constant $C$
  • Near-Optimal Asymptotic Error: $\lVert[L]_{n}\rVert_{2\to\infty}=O(\log^{1+\alpha}(n))$
  • Computability: There exists an unbounded streaming algorithm that at each time step $t$ takes as input $z_{t}$ and outputs $(Lz)_{t}$ using $O(t)$ memory and amortized $O(\log t)$ time
Corollary 1.
For all $\varepsilon,\delta_{priv},\alpha>0$ there exists an unbounded streaming algorithm for continual counting, described in Algorithm 1, which has the same complexity as in Theorem 1 and satisfies $(\varepsilon,\delta_{priv})$-DP in the continual release model. At each time step $t$, the algorithm adds Gaussian noise with scale $O\big(\log^{1+\alpha}(t)C_{\varepsilon,\delta}\big)$, where $C_{\varepsilon,\delta}=O(\varepsilon^{-1}\sqrt{\log(1/\delta_{priv})})$ is independent of the input.
We prove Theorem 1 in section 4; Corollary 1 then follows from existing results.  [li_matrix_nodate, denisov2022improved].
3.3 Technical Overview
The matrix decompositions we consider are of the form $L=\operatorname{LTToep}(f(z;-\gamma,-\delta)),~R=\operatorname{LTToep}(f(z;\gamma,\delta))$, where:
$f(z;\gamma,\delta):=f_{1}(z)f_{2}(z;\gamma)f_{3}(z;\delta)$ (1)
$f_{1}(z):=\frac{1}{\sqrt{1-z}}~,\quad f_{2}(z;\gamma):=\Big(\frac{1}{z}\ln\Big(\frac{1}{1-z}\Big)\Big)^{\gamma}~,\quad f_{3}(z;\delta):=\Big(\frac{2}{z}\ln\Big(\frac{1}{z}\ln\Big(\frac{1}{1-z}\Big)\Big)\Big)^{\delta}$
As motivation for why one might study such matrices, recall that the first column of $M_{\textit{count}}$ corresponds to the Taylor coefficients of $(1-z)^{-1}$. This implies that for any pair of functions $g_{1},g_{2}$ analytic on the open unit disc with product $(1-z)^{-1}$, we have $M_{\textit{count}}=\operatorname{LTToep}(g_{1})\operatorname{LTToep}(g_{2})$. Conversely, any non-pathological decomposition of $M_{\textit{count}}$ into LTToep matrices will correspond to such a pair of functions. So, rather than searching for arbitrary sequences of real numbers (which are complicated to represent and reason about), we can instead take the approach of searching for nice analytic functions with the hope of translating their functional properties into statements about the sequences we’re ultimately interested in.
To understand why these functions in particular are good candidates, consider the special case where $\gamma=\delta=0$. This gives us $R=\operatorname{LTToep}((1-z)^{-1/2})$, which is exactly $M_{\textit{count}}^{1/2}$  [dvijotham_efficient_2024]. The key issue that makes this sequence inapplicable to unbounded inputs is that it is not square-summable. By Parseval’s theorem, this is equivalent to the statement that the function $f$ is not square-integrable. To fix this, we would like to ‘nudge’ $f$ so that it becomes square-integrable while still being as ‘close’ to the original function as possible. A plausible first attempt in this direction would be to choose $R=\operatorname{LTToep}((1-z)^{-1/2+\alpha})$ for some $\alpha>0$, but this gives $O(t^{\alpha})$ error; the key issue is that $(1-z)^{-\alpha}$ diverges very quickly as $z\to 1^{-}$ even when $\alpha$ is small. To obtain logarithmic overhead, we require functions like $\ln(1/(1-z))$ that diverge more slowly. Multiplying by $f_{2}$ turns out to be sufficient to prove Theorem 1, but we show in section 5 that incorporating $f_{3}$ can further improve our variance by a constant factor. Finally, the inclusion of the $1/z$ and $2/z$ terms in Equation 1 eliminates the 0 of $\ln(1/(1-z))$ at $z=0$, which ensures that $f$ is always analytic with $f(0)=1$.
To study the asymptotic growth of $\lVert[L]_{n}\rVert_{2\to\infty}$ and $\lVert[R]_{n}\lVert_{1\to 2}$, we draw on the classic work of flajolet1990singularity, whose Theorem 3B (reproduced fully in Theorem 2) provides the following asymptotic expansion for the coefficient of $z^{n}$ in the Taylor series of $f$:
$z^{n}]f(z;\gamma,\delta)\sim\frac{1}{\sqrt{n\pi}}(\ln n)^{\gamma}(2\ln\ln n)^{\delta}\cdot\Big(1+O(\ln^{-1}(n))\Big)$
In particular, this implies that if $\gamma=-1/2-\alpha$ for some $\alpha>0$, then the series $\sum_{n=0}^{\infty}([z^{n}]f(z;\gamma,\delta))^{2}$ converges, which is equivalent to $R$ having bounded column norm. This is the crucial step that allows us to operate on unbounded time streams without the use of doubling tricks — by calibrating our noise scale to this limit, which can be computed exactly by integrating $|f(z;\gamma,\delta)|^{2}$ over the unit circle using Parseval’s theorem, we can guarantee privacy for all finite inputs. Using the same asymptotic expansion, we can also show that the row norm of the corresponding $L$ matrix grows like $O(\log^{1+\alpha}(n))$ when $\delta\geq 0$ or $O(\log^{1+\alpha+o(1)}(n))$ when $\delta<0$.
Finally, we can compute the first $t$ coefficients of $f_{L}$ and its convolution with $z$ in $O(t\log t)$ time and $O(t)$ space using FFTs. Algorithm 1 combines all of these ideas together alongside a standard doubling trick on $t$, which allows us to achieve $O(n\log n)$ time complexity even when $n$ is unknown.
Algorithm 1 Logarithmic Matrix Method
Matrix parameters $\gamma<-1/2$ and $\delta$, privacy parameter $C_{\varepsilon,\delta}$
Compute $\Delta\leftarrow\lVert\operatorname{LTToep}(f(z;\gamma,\delta)\rVert_{1\to 2}\triangleright$ See ’Bounded Sensitivity’ in section 4
for $t=1,\ldots,n$ do
  Receive input $x_{t}\in[0,1]$
  Set $S_{t}\leftarrow\sum_{i=1}^{t}x_{t}$
  if $t=2^{m}$ for some integer $m$ then
   Sample $z_{s}\sim\mathcal{N}(0,C_{\varepsilon,\delta}^{2}\Delta^{2})$ for $t\leq s\leq 2t-1$
   Compute next $t$ coefficients of $L=\operatorname{LTToep}(f(z;-\gamma,-\delta))$ in $O(t\log t)$ time
   $\triangleright$ See ’Computability’ in section 4
   Compute next $t$ terms of the sequence LzLz in $O(t\log t)$ time with an FFT
  Output $S_{t}+(Lz)_{t}$
4 Proof of Theorem 1
We will begin by fixing $\alpha>0$, and choosing $f_{R}(z)=f(z;-1/2-\alpha,\delta),~f_{L}(z)=f(z;1/2+\alpha,-\delta)$ with $f$ defined as in Equation 1. Our goal is to show that the matrix decomposition $L=\operatorname{LTToep}(f_{L}),R=\operatorname{LTToep}(f_{R})$ satisfies all four properties listed in Theorem 1.
Joint Validity.
We have that $LR=\operatorname{LTToep}(f_{L}f_{R})=\operatorname{LTToep}(\frac{1}{1-z})=M_{\textit{count}}$.
Bounded Sensitivity.
We begin by showing that the column norm of $R$ is finite, or equivalently that $f_{R}\in L^{2}$. Recall that $\lVert R\rVert_{1\to 2}^{2}=\sum_{n=0}^{\infty}([z^{n}]f_{R})^{2}$. By Theorem 3B of flajolet1990singularity, we know that $([z^{n}]f_{R})^{2}=O(n^{-1}\log^{-1-2\alpha}(n)\log^{2\delta}(\log n))=O(n^{-1}\log^{-1-\alpha}(n))$, and the series therefore converges by the integral test.
To actually compute this sensitivity, we make use of Parseval’s theorem, a classic result in Fourier analysis which relates square-summable sequences to functions that are square-integrable over the unit circle. Specifically, it gives us:
$\sum_{m=0}^{\infty}([z^{m}]f_{R})^{2}=\frac{1}{2\pi}\int_{0}^{2\pi}|f_{R}(e^{i\theta})|^{2}~d\theta$
In Appendix A, we simplify the right-hand side into two integrals over smooth, real-valued functions that can be integrated numerically to high precision.
Near-Optimal Asymptotic Error.
We first consider the case when $\delta\geq 0$. Once, again Theorem 3B of flajolet1990singularity tells us that $([z^{n}]f_{L})^{2}=O(n^{-1}(\ln n)^{1+2\alpha}(\ln\ln n)^{-2\delta})=O(n^{-1}(\ln n)^{1+2\alpha})$. So, by the definition of big-$O$, there are constants $n_{0},C$ such that for all $n>n_{0}$, we have $\sum_{m=n_{0}}^{n}([z^{m}]f_{L})^{2}\leq C\sum_{m=n_{0}}^{n}\frac{1}{m}(\ln m)^{1+2\alpha}\leq C\int_{n_{0}-1}^{n}\frac{1}{x}(\ln x)^{1+2\alpha}~dx=\frac{C}{2+2\alpha}(\ln(n)^{2+2\alpha}-\ln(n_{0}-1)^{2+2\alpha})$. Therefore, $\sqrt{\sum_{m=0}^{n}([z^{m}]f_{L})^{2}}=O(\ln(n)^{1+\alpha})$ as desired.
The case where $\delta<0$ is similar, except we have $2\delta>0$ and so it is no longer true that our coefficients are $O(n^{-1}(\ln n)^{1+2\alpha})$. But for any $\alpha^{\prime}>\alpha$, $(\ln\ln n)^{-2\delta}=o((\ln n)^{\alpha^{\prime}-\alpha})$, and so by the same basic argument we can derive an asymptotic error bound of $O(\ln(n)^{1+\alpha+o(1)})$ instead.
Computability
We formalize the problem we are trying to solve as a game: first, an adversary secretly fixes some integer $n>0$. Then, at each time step $t=1,\ldots,n$, our algorithm is required to output $(\operatorname{LTToep}(f_{L})z)_{t}$, where $z\sim\mathcal{N}(0,I_{n})$. The algorithm’s goal is to achieve an optimal asymptotic dependence on the unknown value $n$ with respect to the total memory and computation used. We will model computation in units of $M(t)$, the cost of multiplying two polynomials of degree $t$. Because polynomial multiplication can be implemented through divide-and-conquer algorithms, we assume that for any $k>0$, $M(kt)\sim kM(t)$ asymptotically as $t\to\infty$.
Initially, we disregard the challenge of not having access to $n$ and consider the intermediate problem of computing $\operatorname{LTToep}(f_{L})]_{t}$ for some constant $t$ that we choose ourselves. As a preview of our eventual strategy for handling unbounded inputs, we assume that we have access to a pre-computed solution of size $t/2$. From here, we observe that $\operatorname{LTToep}{f(z;-\gamma,-\delta)}=\operatorname{LTToep}(f_{1}(z))\operatorname{LTToep}(f_{2}(z;\gamma))\operatorname{LTToep}(f_{3}(z;\delta))$. So, it suffices to compute each of these matrices in isolation. The final product can then be computed using $2M(t)$ (or just $M(t)$ if $\delta=0$).
To compute $\operatorname{LTToep}(f_{1}(z))$, we use the fact that $f_{1}=\sum_{m=0}^{\infty}(-x)^{m}{-1/2\choose m}$. This can be used to derive the recurrence relation $z^{0}]f_{1}=1,~[z^{m}]f_{1}=(1-\frac{1}{2m})[z^{m-1}]f_{1}$ presented in prior works  [henzinger_almost_2023, fichtenberger2023constant]. This recurrence lets us compute the coefficients of this matrix in $O(t)$ time and space.
To compute the cofficients of $\operatorname{LTToep}(f_{2}(z;\gamma)$, we begin with the fact that $z^{m}]f_{2}(z;1)=\frac{1}{m+1}$, which can be derived from the Taylor expansion of $\ln(1+z)$. This intermediate sequence can also clearly be computed in linear time and space. To account for the power of $\gamma$, we use the identity $f_{2}(z;-\gamma)=\exp(-\gamma\ln(f_{2}(z;1)))$. To compute $\ln(f_{2}(z;1))$, we use the fact that $\ln^{\prime}(f_{2}(z;1))=\frac{f_{2}^{\prime}(z;1)}{f_{2}(z^{\prime}1)}$. The coefficients of $f_{2}^{\prime}(z;1)$ can be directly computed from the coefficients of $f_{2}(z;1)$ through term-by-term differentiation, and using the division algorithm of hanrot2004newton lets us compute the ratio $f_{2}^{\prime}(z;1)/f_{2}(z)$ using $2.25M(t)$. Term-by-term integration then allows us to recover all but the constant term of $\ln(f_{2}(z;1))$ in linear time, but this is sufficient because we know $\ln(f_{2}(0;1))=\ln(1)=0$.
Having shown that we can efficiently compute the Taylor coefficients of $\gamma\ln(f_{2}(z;1))$, it remains to find the coefficients of $f_{2}(z;-\gamma)=\exp(-\gamma\ln(f_{2}(z;1))$. We achieve this by using the algorithm also presented in  [hanrot2004newton] that takes an order $t/2$ approximation of $\exp(h)$ as input and produces an order $t$ approximation using $2.25M(t)$.
Finally, we can rewrite $f_{3}(z;\delta)$ as $(\frac{2}{z}\ln(f_{2}(z;1))^{\delta}$. The allows us to reuse our earlier computation of the Taylor coefficients of $\ln(f_{2}(z;1))$. Dividing by $z$ corresponds to a simple shift of the coefficients. All that remains is to account for the power of $\delta$, which can be done in $4.5M(t)$ using the same technique described above for $\gamma$. We can improve this to just $2.25M(t)$ if $\delta=-\gamma$ by computing the quotient $f_{2}(z;1)/f_{3}(z;1)$ before applying the power operation. Doing so also saves us an extra $M(t)$ by removing the need to multiply $\operatorname{LTToep}(f_{2}(z;\gamma)$ and $\operatorname{LTToep}(f_{3}(z;\delta)$ at the end.
In total, we require $11M(t)$ to compute $L]_{t}$, plus one more $M(t)$ to compute $L]_{t}z$. To extend to the case of general $n$, we initialize tt=2 and double it when we are asked to output the t+1t+1st estimate. This gives us a total cost over the entire input of $12\sum_{k=1}^{\lceil\log_{2}n\rceil}M(2^{k})\sim 12\sum_{k=1}^{\lceil\log_{2}n\rceil}2^{-k+1}M(n)\leq 24M(n)$, which can be cut down to $17.5M(n)$ if $\delta=-\gamma$, or $13M(n)$ when $\delta=0$. When polynomial multiplication is implemented with FFTs, we arrive at $O(n\log n)$ time and $O(n)$ space complexity, asymptotically matching the algorithm of henzinger_almost_2023. But note that because we have to adjust $t$ on the fly, our algorithm requires amortized $O(\log t)$ time at each step, compared to $O(n\log n)$ pre-processing and $O(1)$ time per step for henzinger_almost_2023.
5 Implementation and Extensions
In this section, we investigate various practical considerations in the implementation of Algorithm 1. We also consider relaxing the assumption that $n$ is completely unknown by giving the algorithm access to (possibly unreliable) side information on its value, and propose an approximate variant (Algorithm 2) which sacrifices some theoretical rigor for a greater than $5\times$ improvement in running time for large inputs. Finally, we show that Algorithm 1 can be used as a subcomponent of the hybrid mechanism of chan2011private to achieve exactly $O(\log^{2}t)$ variance with improved constants compared to their original approach.  [^1] Our code is available at https://github.com/ben-jacobsen/central-dpolo/
Choosing the value of $\alpha$.
While it is in principle possible to set $\alpha$ arbitrarily small, this would be unwise. To see why, observe that for the purposes of bounding sensitivity, there is a stark qualitative gap between the matrix $\operatorname{LTToep}(f(z;-1/2-10^{-100},\delta))$, which has finite column norm, and the matrix $\operatorname{LTToep}(f(z;-1/2,\delta))$ whose column norm diverges. But at the same time, the row norms of the matrices $\operatorname{LTToep}(f(z;1/2+10^{-100},-\delta))$ and $\operatorname{LTToep}(f(z;1/2,-\delta))$ both diverge at pretty much the same rate. This asymmetry implies that reducing $\alpha$ past a certain point does little to improve our actual error while inviting numerical instability.
So, how should $\alpha$ be chosen? We observe that, as of mid-2025, the world’s population is estimated to be about 8.2e9. If every single one of those people contributed 100 data points to our algorithm, then we would have $\ln^{0.01}(n)\approx 1.03$. We conclude that setting $\alpha=0.01$ is likely a safe and practical choice for most applications.
Choosing the value of $\delta$.
[…]
Exploiting imperfect information about $n$.
Up to this point we have assumed that $n$ is unknown and impossible to predict, but in practice, history or expert knowledge might suggest a range of plausible values. We model this side information with the double-inequality $n_{0}\leq n\leq Cn_{0}$ for some $n_{0},C>0$, which is given to the algorithm. We allow these bounds to be unreliable in two senses — the upper bound might be loose, or it might be entirely wrong. We show that our algorithm can make efficient use of this information when it is reliable without incurring any additional error when it isn’t.
Our basic strategy is to simply pre-compute $Lz$ out to $Cn_{0}$ terms. Provided that the upper bound is true, we recover the same $O(n\log n)$ pre-processing time and $O(1)$ work per-iteration as the $M_{\textit{count}}^{1/2}$ algorithm. Moreover, we suffer no performance cost if the upper bound turns out to be loose because our error at each round depends only on $t$. In contrast, if we used the $M_{\textit{count}}^{1/2}$ algorithm with input $Cn_{0}$, we would be forced to calibrate its noise scale to the sensitivity of the larger matrix $M_{\textit{count}}^{1/2}]_{Cn_{0}}$. In the unlucky event where it turns out that $n=n_{0}$, this would produce a uniform $\sim 1+\log C/(1+\log n_{0})$ multiplicative scaling of variance across the entire input.
In the other direction, if the upper bound turns out to be false then our algorithm can recover gracefully using the same doubling trick described in section 4. But by that point, the $M_{\textit{count}}^{1/2}$ algorithm will have completely exhausted the privacy budget of the first data-point, forcing it to either weaken its privacy guarantees post-hoc or restart entirely with a new estimate for the input size! Even if this new estimate is perfectly correct, the initial mistake will still lead to $O(\log(Cn_{0}))$ additive error.
Constant-factor speedups through asymptotic expansions.
flajolet1990singularity derive the following asymptotic expansion for the Taylor coefficients of $f$, which we used earlier in a weaker form to derive the asymptotic error rate of our mechanism:
Theorem 2 (Paraphrased from flajolet1990singularity Theorem 3B).
Let $\gamma,$ and $\delta$ be complex numbers not in $\{0,1,2,\ldots\}$. Then the Taylor coefficients of $f(z;\gamma,\delta)$ satisfy:
$z^{m}]f(z;\gamma,\delta)\sim\frac{1}{\sqrt{m\pi}}(\ln m)^{\gamma}(\ln\ln m)^{\delta}\Big(1+\sum_{k\geq 1}\frac{e_{k}^{(\gamma,\delta)}(\ln\ln m)}{(\ln(m)\ln(\ln(m)))^{k}}\Big)$ (2)
where $e_{k}^{(\gamma,\delta)}(x)$ is a polynomial of degree $k$:
$e_{k}^{(\gamma,\delta)}(x)=\sqrt{\pi}\frac{d^{k}}{ds^{k}}\big(\frac{1}{\Gamma(-s)}\big)\big|_{s=-1/2}\cdot E_{k}(x)$
and $E_{k}(x)$ is the $k^{th}$ Taylor coefficient of the function $g(u)=(1-xu)^{\gamma}\big(1-\frac{1}{x}\ln(1-xu)\big)^{\delta}$.
While this expression is messy, the cost of computing it is independent of $m$: for fixed approximation order $K$, the derivatives of the reciprocal gamma function can be precomputed and the remaining terms can be calculated in $O(K\log K)$ time using the same techniques described in section 4. This idea naturally suggests Algorithm 2, which switches from exact computations to order $K$ approximations once the relative error of the approximation falls below some threshold $\eta$. We plot relative error as a function of $t,\delta$ and $K$ in Figure 1.
Figure 1: Plots the relative error $|\hat{r}_{t}-r_{t}|/r_{t}$ from Algorithm 2 as a function of $t$ when $\gamma=-0.51$. The asymptotic expansion converges much more quickly when $\delta=0$.
We highlight a potential pitfall with this approach, which is that if we approximate $\hat{L}\approx L$ directly, then we are implicitly choosing $\hat{R}=M_{\textit{count}}\hat{L}^{-1}$ and can no longer calibrate our noise to $\lVert R\rVert_{1\to 2}$ as normal. Ideally we would like to be able to prove a bound like $\lVert M_{\textit{count}}\hat{L}^{-1}\rVert_{1\to 2}\leq(1+O(\eta))\lVert R\lVert_{1\to 2}$, but this appears to be non-trivial. We therefore take the opposite approach, which is to directly approximate $\hat{R}\approx R$ and calibrate our noise to $(1+\eta)\lVert R\rVert_{1\to 2}$ with $\hat{L}=\hat{R}^{-1}M_{\textit{count}}$. This version of the algorithm is provably valid and private, and for large enough inputs it reduces the computational cost at each power of 2 from $12M(t)$ to $2.25M(t)$ (the cost of one series division), closing over $80\%$ of the computational gap between our algorithm and a standard doubling trick. The tradeoff is a multiplicative $(1+\eta)$ increase in error prior to the switch, and the loss of tight, provable bounds on error after the switch. In Figure 1, we compare the performance of the two algorithms, and find that after switching around $t=2^{11}$, the performance of Algorithm 2 remains indistinguishable from Algorithm 1 out to $t=2^{24}$.
Algorithm 2 Approximate Logarithmic Matrix Method
Matrix parameters $\gamma,\delta\not\in\{0,1,\ldots\}$, order $K$, error tolerance $\eta$, privacy parameter $C_{\varepsilon,\delta}$
for $t=1,\ldots,n$ do
  if $t=2^{m}$ for some integer $m$ then
   Sample $z_{s}\sim\mathcal{N}(0,(1+\eta)^{2}C_{\varepsilon,\delta}^{2})$ for $t\leq s\leq 2t-1$
   $r_{t}\leftarrow[z^{t-1}]f_{R}$ using section 4 (’Computability’)
   $\hat{r}_{t}\leftarrow$ Degree $K$ approximation of $z^{t-1}]f_{R}$ from Equation 2
   if $\frac{|r_{t}-\hat{r}_{t}|}{r_{t}}>\eta$ then
     Use section 4 to compute next $t$ entries of $\operatorname{LTToep}(f_{R})$
   else
     Use Equation 2 to approximate the next $t$ entries of $\operatorname{LTToep}(f_{R})$
   Save computed sequence as $\hat{R}]_{2t}$
   Compute $\hat{R}]_{2t}^{-1}[M_{\textit{count}}]_{2t}z$
  Output $(M_{\textit{count}}\hat{R}^{-1}z)_{t}$
Asymptotic improvements through hybrid mechanisms.
In their seminal paper, chan2011private propose a generic ‘Hybrid Mechanism’ for continual counting of unbounded streams. The schema requires one (possibly inefficient) unbounded counting algorithm, which operates on the condensed sequence $y_{k}=\sum_{t=2^{k}}^{2^{k+1}-1}x_{t}$, and a bounded algorithm, which is restarted whenever $t$ is a power of 2. The privacy budget is divided between these two algorithms, and the partial sum at time $t=2^{k}+r$ is given by adding together estimates for $\sum_{i=0}^{k-1}y_{i}$ and $\sum_{t=2^{k}}^{2^{k}+r}x_{t}$. The resulting hybrid mechanism is unbounded and typically preserves the exact asymptotic error rate of the chosen bounded algorithm.
The original hybrid mechanism used a simple unbounded algorithm that adds independent noise to each input (corresponding to the matrix decomposition $L=M_{\textit{count}},R=I$). Our work enables a more powerful instantiation using Algorithm 1 as the unbounded algorithm instead (‘Hybrid Matrix’ in Table 1), which we compare against the original. For both implementations, we improve on the original presentation by using $M_{\textit{count}}^{1/2}$ for the bounded algorithm and allocating 75% of the privacy budget to the bounded learner, which reduces the constant of the leading $O(\log^{2}t)$ term. In the spirit of honaker2015efficient, we also reuse the outputs of the bounded mechanism on earlier subsequences to optimize the variance of the final estimate. Our results are visualized in Figure 2.
Figure 2: Left: Comparison of the exact variance of different algorithms and parameter choices. Contrast with asymptotics in Table 1. The Hybrid mechanism using Algorithm 1 for the unbounded component outperforms the variant using independent noise as in chan2011private, but both variants exhibit very erratic performance when $t$ is close to a power of 2. For Algorithm 1, the parameters $\delta=-\gamma$ and $\delta=-6\gamma/5$ give similar performance, but the latter is slightly better once $t>2^{20}$, and both outperform $\delta=0$ when $t>2^{10}$. Note that the performance of Algorithm 2 with $\delta=0$ is visually indistinguishable from that of Algorithm 1. Right: Coefficients of the matrices $\operatorname{LTToep}(f_{L})$ (upper) and $\operatorname{LTToep}(f_{R})$ (lower) for various choices of $\delta$. The choice $\delta=-6\gamma/5$ produces lines that are as close as possible to that of $M_{\textit{count}}^{1/2}$ without crossing it, which we hypothesize explains its good performance.
6 Conclusion
We have shown for the first time that the classic matrix factorization method [li_matrix_nodate] can be efficiently extended to online settings with unknown input size without the use of doubling tricks. The resulting algorithm is the first that we are aware of that is simultaneously unbounded and smooth with almost-optimal asymptotic error. Empirically, we have also shown that it enjoys excellent constants when its parameters are set correctly, with variance that is uniformly less than $1.5\times$ that of henzinger_almost_2023 for inputs as large as $n=2^{24}$.
Many interesting questions remain unanswered, however. For instance: the only algorithms we are aware of that exactly achieve asymptotically optimal $O(\log^{2}t)$ variance in the unbounded setting rely on some form of doubling trick, and despite a great deal of effort, we were not able to find a $\operatorname{LTToep}$ decomposition with this property. Our suspicion is that no such decomposition exists, and that this is related to the classic result that the sequence $a_{n}=(n\cdot\log n\cdot\log\log n\cdot\ldots\cdot\log^{\circ k}n)^{-1}$ diverges for any $k$. We believe that a formal proof of this conjecture would be very interesting as it would constitute a clean separation between the power of $\operatorname{LTToep}$ and general lower-triangular matrices in online learning. We therefore conclude by posing the following open question:
Open Question 1.
Do there exist $\operatorname{LTToep}$ matrices $L,R$ such that $L]_{n}[R]_{n}=[M_{\textit{count}}]_{n}$, $\lim_{n\to\infty}\lVert[R]_{n}\rVert_{1\to 2}<\infty$, and $\lVert[L]_{n}\rVert_{2\to\infty}=\Theta(\log^{2}n)$?
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction are exactly the content of Theorem 1, which we prove.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Our work has two main limitations: the computational cost of computing the matrices we consider, and the slightly-suboptimal asymptotic error growth. These issues are discussed in detail in section 4 and the conclusion, respectively.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Our primary theoretical result is Theorem 1, which is proved fully in section 4. Other minor results in section 5 are presented with justification. One computational proof, related to the sensitivity calculation, is deferred to the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our experiments all relate to non-random numerical quantities that are explicitly defined in the paper, and we offer a very thorough description of the algorithm used to compute them in section 4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have released our code as open-source software, including scripts to reproduce all figures in the paper. It is available at https://github.com/ben-jacobsen/central-dpolo/
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We only conduct two small experiments, which are described in Figure 1 and Figure 2. We provide all parameters for our own algorithm, and explicitly describe the construction of the hybrid mechanisms. The Smooth Binary and Sqrt Matrix algorithms have no hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: Our experiments are all exact and involve no randomness.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No]
  39. Justification: There are no special computational resources required to reproduce our experiments. All of them were carried out in python using a single commercial CPU in a few seconds or minutes.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work is largely theoretical and defensive in nature, and does not involve any human subjects or sensitive datasets.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: The work is largely theoretical and not tied to any concrete application. We do not anticipate any particular negative societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper does not release any data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We do not use any existing data or models. The primary package we use is mpmath, which we cite in Appendix A.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper did not use LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Technical Details for Computing Sensitivity
To actually compute the sensitivity, we make use of Parseval’s theorem, which gives us:
$\sum_{n=0}^{\infty}([z^{n}]f_{R})^{2}=\frac{1}{2\pi}\int_{0}^{2\pi}\lvert f_{R}(e^{i\theta})\rvert^{2}~d\theta$
$=\frac{2^{2\delta}}{2\pi}\int_{0}^{2\pi}(2\sin\Big(\frac{\theta}{2}\Big))^{-1}\Big(\ln^{2}(2\sin\Big(\frac{\theta}{2}\Big))+\frac{(\pi-\theta)^{2}}{4}\Big)^{-1/2-\alpha}\Bigg[\frac{1}{4}\ln^{2}\Big(\ln^{2}(2\sin\Big(\frac{\theta}{2}\Big))+\frac{(\pi-\theta)^{2})}{4}\Big)$
$\qquad\qquad+\Big(\operatorname{atan2}2\Big(\frac{\pi-\theta}{2},-\ln\big(2\sin\Big(\frac{\theta}{2}\Big)\big)\Big)+\operatorname{atan2}2(-\sin\theta,\cos\theta)\Big)^{2}~\Bigg]^{\delta}~d\theta$
Where $\operatorname{atan2}2(y,x)=\arctan(y/x)$, possibly shifted by $\pm\pi$ to fall within $(-\pi,\pi]$. With the change of variables $\omega=(\pi-\theta)/2$, this simplifies to:
$\frac{2^{2\delta}}{\pi}\int_{-\pi/2}^{\pi/2}$ $(2\cos\omega)^{-1}\Big(\ln^{2}(2\cos\omega)+\omega^{2}\Big)^{-1/2-\alpha}\Bigg[\frac{1}{4}\ln^{2}\Big(\ln^{2}(2\cos\omega)+\omega^{2}\Big)$
$+\Big(\operatorname{atan2}2(\omega,-\ln(2\cos w))+\operatorname{atan2}2(-\sin(2\omega),-\cos(2\omega))\Big)^{2}~\Bigg]^{\delta}~d\omega$
From here, we define the functions:
$I_{1}(\omega)=2\cos\omega\quad I_{2}(\omega)=\ln^{2}(I_{1}(\omega))+\omega^{2}\quad I_{3}(\omega)=\frac{1}{4}\ln^{2}(I_{2}(\omega))\quad I_{4}(\omega)=\arctan\big(-\omega/\ln(I_{1}(\omega))\big)+2\omega\big)$
By using symmetry and splitting the interval based on the offset of $\operatorname{atan2}2$, we finally arrive at:
$\lVert R\rVert_{1\to 2}^{2}=\frac{2^{1+2\delta}}{\pi}\Bigg[$ $\int_{0}^{\pi/3}I_{1}(\omega)^{-1}I_{2}(\omega)^{-1/2-\alpha}(I_{3}(\omega)+I_{4}(\omega)^{2})^{\delta}~d\omega$
$+$ $\int_{\pi/3}^{\pi/2}I_{1}(\omega)^{-1}I_{2}(\omega)^{-1/2-\alpha}(I_{3}(\omega)+(I_{4}(\omega)-\pi)^{2})^{\delta}~d\omega~\Bigg]$
Expressed in this way, both integrands are sufficiently smooth to be numerically integrated to high precision. We use mpmath [johansson2013mpmath] for this purpose in our experiments, which supports arbitrary-precision floating point computations and numerical integration.